Consider this reaction: H 2 (g)+ I 2 (g)⇌2HI(g) H2(g)+I2(g)⇌2HI(g) A rxn mixture in a 3.73 L flask at a certain temperature initially contains 0.766 g H2 and 96.7 g I 2 I2 . At equilibrium, the flask holds 90.3 g HI .

Calculate the equilibrium constant (Kc)(Kc) for the reaction at this temperature.

Record your answer using two significant figures.
Answer #1

Salution- Holg) + I2Cg) - 2 HIG) 0.766 m = 0-381720le juitel moles of m) = 2.019u initial males of initial males of 29 Eg 96.DATE / HIJ 2 2х - усце 2 2Х o-3524 3-73 [н]> Кс 2 . на] , ko - ©- 1892) 7:56Xic® X251x46 ke = 3-5797x10-2 5°C8 x1-5 Ке6: 32Х